OpenStudy (anonymous):

the sequence:5,25,125,625 what is the 7th term of the sequence. what algebric expression represents the nth term

OpenStudy (anonymous):

the 7th term is 78125

OpenStudy (anonymous):

@sweetsunray

OpenStudy (anonymous):

\[5 \times 5^6\]

OpenStudy (anonymous):

Well, what pattern did you use to find the 7th term? Replace 7 with n to find an algebraic representation of the nth term.

OpenStudy (anonymous):

i multply by 5

OpenStudy (anonymous):

this is a geometric sequence @NinjaDevo

OpenStudy (anonymous):

a1 * 5^(7-1)

OpenStudy (anonymous):

That's what I was getting at...just replace the 7 with a n....

OpenStudy (anonymous):

@NinjaDevo when i solve that eqaution with a1 * 5^(7-1) it doesnt give me 78125 which is the 7th term

OpenStudy (anonymous):

@math&ing001

OpenStudy (anonymous):

@phi

OpenStudy (math&ing001):

@Pandabear78963 your answer is correct.

OpenStudy (anonymous):

i know but what will the equation be to represent the nth term @math&ing001

OpenStudy (phi):

5,25,125,625 is 5^1, 5^2, 5^3, 5^4 it looks like 5^n where n is the term number 5^7 is the answer. we should use a calculator

OpenStudy (anonymous):

78125

OpenStudy (math&ing001):

As @phi said 5^n.

OpenStudy (phi):

***when i solve that eqaution with a1 * 5^(7-1) **** it should: 5* 5^6 = 5^7

OpenStudy (anonymous):

thsnks @phi @math&ing001

OpenStudy (math&ing001):

Welcome !